# Ultrastable jammed sphere packings with a wide range of particle dispersities

### v1

@@ IV Discussion and Conclusions
Many real granular materials are highly polydisperse. For example, power-law-like particle-size distributions $P(\sigma)\sim\sigma^{-\beta}$ over two or more orders of magnitude in $\sigma$ are found in natural systems ranging from soils Wu et al. (1993) to sea ice Rothrock and Thorndike (1984), and in man-made materials produced by a wide range of comminution/fragmentation processes Turcotte (1986). Simulations can systematically study the effects of this dispersity by employing variable-width $P(\sigma)$ with $\mathcal{R}$-dependent $\Delta$, e.g. those given by Eqs. 1-2, but suprisingly few studies have done so. In particular, the preparation-protocol-dependence (PPD) of amorphous solids’ structure and mechanics remains a very active field of research with many open questions Jin et al. (2018); et al. (2020); Berthier et al. (2025), but most theoretical studies of these questions have employed models with a single, fixed particle dispersity $\Delta\lesssim 0.25$.
@@
In this paper, we have shown that the dramatic increases in jammed sphere packings’ ($\phi_{\rm J},\Psi$) achievable by thermally equilibrating them at a higher packing fraction $\phi_{\rm eq}$ Chaudhuri et al. (2010); Charbonneau and Morse (2021); Berthier et al. (2016a); Ozawa et al. (2017), compressing them slower Torquato et al. (2000), artificially speeding up their dynamics during sample compression Ghimenti et al. (2024); Berthier et al. (2024), or introducing additional degrees of freedom associated with the particle diameters Kapteijns et al. (2019); Hagh et al. (2022); bol can be achieved for a very wide range of particle dispersities ($0<\Delta\leq 0.50$), but the effects of employing such techniques can vary strongly and nontrivially with $\Delta$.
@@
We showed that the packing-efficiency/mechanical stability gains obtained by employing SWAP and TDOF moves increase monotonically with $\Delta$. This result was surprising because the utility of standard SWAP for equilibrating deeply supercooled liquids decreases for $\Delta\gtrsim 0.25$ Ninarello et al. (2017); we believe that it holds because our algorithm still yields $\gg 1$ successful swaps per particle over the course of the particle-growth process. As summarized in Section III, packing-efficiency/mechanical stability gains tend to increase rapidly over the range $0<\Delta\lesssim 0.1$, then more slowly over the range $0.1<\Delta\lesssim 0.25$, then linearly over the range $0.25<\Delta\lesssim 0.45$, then more-rapidly for even larger $\Delta$ as the effects of including TDOF moves at the end of the process grow stronger.
@@
Including the latter moves offers several advantages over algorithms that only employ SWAP. For example, they allow one to achieve both higher $\phi_{\rm J}$ and lower $f_{\rm ratt}$ Hagh et al. (2022); bol , whereas standard algorithms yield increasing $f_{\rm ratt}(\phi_{\rm J})$ Ozawa et al. (2017); Monti et al. (2022, 2023). This combination allows one to produce packings that are unusually mechanically/vibrationally stable at $T=0$, and have ideal-glass-like properties at finite $T$ bol . We have shown that they can double the packings’ bulk moduli for $\Delta\gtrsim 0.45$, and attributed these large gains to the fact that TDOF moves allow small rattlers that are not initially part of the packings’ mechanically-rigid backbones to continue growing until until they, too, jam. Even-more-dramatic stability gains are expected for the packings’ shear moduli $G(\Delta)$ because suppression of nonaffine deformation increases $G$ much more than it increases $K$ O’Hern et al. (2003).
@@
Here we only studied systems with $\beta=3$; this value was chosen primarily because it has been employed in multiple recent studies of the glass-jamming transition Interiano-Alberto et al. (2024); Ozawa et al. (2017); Berthier et al. (2024, 2016b); Ninarello et al. (2017); Scalliet et al. (2022); Monti et al. (2022, 2023). We anticipate, however, that the qualitative trends discussed above will also be present for other values of $\beta$. $\beta=3.5$ should be an especially-interesting case for further study, as it yielded the highest $\phi_{\rm J}(\Delta)$ in Refs. Monti et al. (2022, 2023), and also corresponds to the $P(\sigma)$ of grains generated by explosive rock fragmentation Turcotte (1986).
We also anticipate that at least some of these trends will be present for other short-ranged purely-repulsive interaction potentials $U(r)$. If so, this will present multiple opportunities to answer open questions, because the differences between ultrastable/ideal glasses and their less-stable counterparts are comparable to the differences between high-($\phi_{\rm J},\Psi$) jammed states and their lower-($\phi_{\rm J},\Psi$) counterparts Swallen et al. (2007); Ediger (2017); Kapteijns et al. (2019); Hagh et al. (2022); bol ; Kim and Hilgenfeldt (2024); Berthier et al. (2025). Exceptionally-stable soft-sphere glasses might be produced by equilibrating systems at high $\phi_{\rm eq}$ using thermal cSWAP Nishikawa et al. (2025), continuing to perform these cSWAP moves during subsequent particle growth/sample compression to $\phi>\phi_{\rm J}$, and then performing TDOF moves comparable to those discussed above at the end of the packing-generation procedure.
@@
This material is based upon work supported by the National Science Foundation under Grant Nos. DMR-2026271 and DMR-2419261.

### v2

@@ IV Discussion and Conclusions
Many real granular materials are highly polydisperse. For example, power-law-like particle-size distributions $P(\sigma)\sim\sigma^{-\beta}$ over two or more orders of magnitude in $\sigma$ are found in natural systems ranging from soils Wu et al. (1993) to sea ice Rothrock and Thorndike (1984), and in man-made materials produced by a wide range of comminution/fragmentation processes Turcotte (1986). Simulations can systematically study the effects of this dispersity by employing variable-width $P_{{\color[rgb]{0,0.88,0}\mathcal{R}}}(\sigma)$ with $\mathcal{R}$-dependent $\Delta$, e.g. those given by Eqs. 1-2, but suprisingly few studies have done so. In particular, the preparation-protocol-dependence of amorphous solids’ structure and mechanics remains a very active field of research with many open questions Jin et al. (2018); et al. (2020); Berthier et al. (2025), but most theoretical studies of these questions have employed models with a single, fixed particle dispersity $\Delta\lesssim 0.25$.
@@
In this paper, we have shown that the dramatic increases in jammed sphere packings’ ($\phi_{\rm J},\Psi$) achievable by thermally equilibrating them at a higher packing fraction $\phi_{\rm eq}$ Chaudhuri et al. (2010); Charbonneau and Morse (2021); Berthier et al. (2016a); Ozawa et al. (2017), compressing them slower Torquato et al. (2000), artificially speeding up their dynamics during sample compression Ghimenti et al. (2024); Berthier et al. (2024), or introducing additional degrees of freedom associated with the particle diameters Kapteijns et al. (2019); Hagh et al. (2022); bol can be achieved for a very wide range of particle dispersities ($0<\Delta\leq 0.50$), but the effects of employing such techniques can vary strongly and nontrivially with $\Delta$. The $\Psi$ investigated here included $Z_{\rm J}$, $P(Z_{i})$, $g(r)$ for intermediate $r$, and $S(q)$ for intermediate $q$; we anticipate that comparable effects will be observable for other $\Psi$.
@@
We showed that the packing-efficiency/mechanical stability gains obtained by employing SWAP and TDOF moves increase almost monotonically with $\Delta$. This result was surprising because the utility of standard SWAP for equilibrating deeply supercooled liquids decreases for $\Delta\gtrsim 0.25$ Ninarello et al. (2017); we believe that it holds because our algorithm still yields $\gg 1$ successful swaps per particle over the course of the particle-growth process. As summarized in Section III, packing-efficiency/mechanical stability gains tend to increase rapidly over the range $0<\Delta\lesssim 0.1$, then slowly over the range $0.1<\Delta\lesssim 0.25$, then linearly over the range $0.25<\Delta\lesssim 0.45$, then more-rapidly for $\Delta\gtrsim 0.45$ as the effects of including TDOF moves at the end of the process grow stronger.
@@
Including the latter moves offers several advantages over algorithms that only employ SWAP. For example, they allow one to achieve both higher $\phi_{\rm J}$ and lower $f_{\rm ratt}$ for $\Delta\gtrsim 0.25$ Hagh et al. (2022); bol ; Kim and Hilgenfeldt (2024), whereas standard algorithms yield increasing $f_{\rm ratt}(\phi_{\rm J})$ Ozawa et al. (2017); Monti et al. (2022, 2023). This combination has been shown to produce packings that are unusually mechanically/vibrationally stable at $T=0$, and have ideal-glass-like properties at finite $T$ bol . We have shown that it can nearly double the packings’ bulk moduli for $\Delta\gtrsim 0.45$, and attributed these large gains to the fact that TDOF moves allow small rattlers that are not initially part of the packings’ mechanically-rigid backbones to continue growing until they, too, jam. Even-more-dramatic stability gains are expected for the packings’ shear moduli $G(\Delta)$ because the suppression of nonaffine deformation associated with the increased $Z_{\rm J}$ should increase $G$ much more than it increases $K$ O’Hern et al. (2003); Zaccone and Scossa-Romano (2011).
@@
Here we only studied systems with $\beta=3$; this value was chosen primarily because it has been employed in multiple recent studies of the glass-jamming transition Interiano-Alberto et al. (2024); Ozawa et al. (2017); Berthier et al. (2024, 2016b); Ninarello et al. (2017); Scalliet et al. (2022); Oquendo-Patino and Estrada (2020, 2022); Monti et al. (2022, 2023); Anzivino et al. (2023); Srivastava et al. (2025). We anticipate, however, that the qualitative trends discussed above will also be present for other values of $\beta$. $\beta=3.5$ should be an especially-interesting case for further study, as it yielded the highest $\phi_{\rm J}(\Delta)$ in Refs. Oquendo-Patino and Estrada (2020, 2022); Monti et al. (2022, 2023), and also corresponds to the $P(\sigma)$ of grains generated by explosive rock fragmentation Turcotte (1986). We also anticipate that at least some of these trends will be present for other short-ranged purely-repulsive interaction potentials $U(r)$. If so, this will present multiple opportunities to answer open questions, because the differences between ultrastable/ideal glasses and their less-stable counterparts are comparable to the differences between high-($\phi_{\rm J},\Psi$) jammed states and their lower-($\phi_{\rm J},\Psi$) counterparts Swallen et al. (2007); Ediger (2017); Kapteijns et al. (2019); Hagh et al. (2022); bol ; Kim and Hilgenfeldt (2024); Berthier et al. (2025). Exceptionally-stable soft-sphere glasses have been prepared recently using thermal cSWAP Nishikawa et al. (2025). Adding TDOF moves to the end of such sample-preparation procedures could make them even more stable.
@@
We thank Joseph Monti and Alessio Zaccone for helpful discussions. This material is based upon work supported by the National Science Foundation under Grant Nos. DMR-2026271 and DMR-2419261.
Appendix A Comparison to Eq. 5 and its associated theory
Anzivino et al. obtained both Eq. 5 and analytic expressions for its parameters $\{a_{1},a_{2},a_{3},b_{1},b_{2},b_{3}\}$ Anzivino et al. (2023):
$\begin{array}[]{c}a_{1}=\displaystyle\frac{5-(8C_{0}+5)\phi_{\rm mono}^{\rm RCP}}{2C_{0}(1-\phi_{\rm mono}^{\rm RCP})Z^{\prime}(\phi_{\rm mono}^{\rm RCP})}\\ \\ a_{2}=\displaystyle\frac{6[1-(C_{0}+1)\phi_{\rm mono}^{\rm RCP}]}{C_{0}(1-\phi_{\rm mono}^{\rm RCP})Z^{\prime}(\phi_{\rm mono}^{\rm RCP})}\\ \\ a_{3}=\displaystyle\frac{(4C_{0}+1)\phi_{\rm mono}^{\rm RCP}}{2C_{0}(1-\phi_{\rm mono}^{\rm RCP})Z^{\prime}(\phi_{\rm mono}^{\rm RCP})}\\ \\ b_{1}=\displaystyle\frac{7}{2}+\displaystyle\frac{4}{(1-\phi_{\rm mono}^{\rm RCP})^{2}Z^{\prime}(\phi_{\rm mono}^{\rm RCP})}\\ \\ b_{2}=3+\displaystyle\frac{6}{(1-\phi_{\rm mono}^{\rm RCP})^{2}Z^{\prime}(\phi_{\rm mono}^{\rm RCP})}\\ \\ b_{3}=\displaystyle\frac{1}{2}-\displaystyle\frac{2}{(1-\phi_{\rm mono}^{\rm RCP})^{2}Z^{\prime}(\phi_{\rm mono}^{\rm RCP})}\\ \end{array}.$ (9)
Here $\phi_{\rm mono}^{\rm RCP}\equiv\phi_{\rm J}(0)$, $\mathcal{Z}=\frac{P}{\rho k_{B}T}$ is the compressibility factor, and $\mathcal{Z}^{\prime}(x)=\frac{d\mathcal{Z}}{d\phi}|_{x}$ can be evaluated using a liquid-state-theoretic equation of state for $\mathcal{Z}(\phi)$. We choose to employ the Carnahan-Starling expression Carnahan and Starling (1969)
$\mathcal{Z}_{\rm CS}(\phi)=\displaystyle\frac{1}{1-\phi}+\displaystyle\frac{3\phi}{(1-\phi)^{2}}+\displaystyle\frac{\phi^{2}(3-\phi)}{(1-\phi)^{3}}.$ (10)
Thus Eq. 5 is a function only of $\phi_{\rm mono}^{\rm RCP}$, $C_{0}$ and $\Delta$.
Figure 8: Comparison of jamming densities for $N=1000$ to the theory of Anzivino et al. The dotted orange and purple curves respectively show the predictions of Eq.  5 for fcc and bcc boundary conditions Anzivino et al. (2023), i.e. for ($\phi_{\rm mono}^{\rm RCP}=0.6774$, $C_{0}={\color[rgb]{0,0,1}0}.01874$) and ($\phi_{\rm mono}^{\rm RCP}=0.6506$, $C_{0}={\color[rgb]{0,0,1}0}.0246$). The green dashed curve show the predictions of Eq.  5 for our measured $\phi_{\rm mono}^{\rm RCP}=0.646$ and fit $C_{0}=.02118$, and inset shows the fractional error of this prediction relative to our simulation data for $\phi_{\rm J}(\Delta)$.
Anzivino et al. chose to calculate both $C_{0}$ and $\phi_{\rm mono}^{\rm RCP}$ from first principles. Their calculation required assuming that the liquid will form either a fcc crystal or a bcc crystal at sufficiently large $\phi$; these two different “boundary conditions” yielded different values of $\phi_{\rm mono}^{\rm RCP}$ and $C_{0}$ Anzivino et al. (2023). Figure 8 compares the predictions of their theory to our data for $\phi_{\rm J}(\Delta)$ in systems prepared without SWAP or TDOF moves. Both predictions qualitatively capture the observed trends, but exhibit substantial quantitative disagreement with our simulation data; Ref. Anzivino et al. (2023) reported similar quantitative disagreements.
An alternative approach is to take the simulated $\phi_{\rm mono}^{\rm RCP}\equiv\phi_{\rm J}(0)$ as a known input, and then fit the data for $\phi_{\rm J}(\Delta)$ to Eq. 5 with $C_{0}$ as the only fitting parameter. Using this approach, we obtain $C_{0}\simeq 0.0212$. As shown in Fig. 8, this value predicts the simulated $\phi_{\rm J}(\Delta)$ to within less than 0.3% over the entire range $0\leq\Delta\leq 0.5$. This agreement is remarkable given that Eq. 9 takes only quantities associated with monodisperse systems [i.e., $\phi_{\rm mono}^{\rm RCP}$ and $Z^{\prime}(\phi_{\rm mono}^{\rm RCP})$] as its inputs and Eq. 10 is an equation of state for monodisperse spheres.
On the other hand, Eq.  5 clearly does not accurately predict even the shape of the $\phi_{\rm J}(\Delta)$ curve for systems prepared using SWAP and TDOF moves. A similar issue arises for ellipses, where a physically-motivated analytic functional form (Eq. 13 of Ref. Rocks and Hoy (2023)) predicts the $\phi_{\rm J}(\alpha)$ of ellipse packings to within less than 0.3% over the range $1\leq\alpha\leq 10$ for systems prepared without these moves, but fails to capture the shape of $\phi_{\rm J}(\alpha)$ for systems prepared with these moves Hoy (2024). These results indicate that the preparation-protocol-dependence of $\phi_{\rm J}$ couples to $\Delta$ and $\alpha$ in a way that is not accounted for by existing theories. Resolving this issue could be a critical contribution to the ongoing effort to analytically predict the structural and mechanical properties of jammed packings Zaccone (2022, 2025a, 2025b).
Appendix B Rattler-related considerations
In soft-sphere packings, contacts are unambiguously defined as interparticle overlaps: $Z_{i}=\sum_{j\neq i}\Theta(\sigma_{ij}-r_{ij})$ where $\Theta$ is the Heaviside step function, and the average coordination number $Z=N^{-1}\sum_{i=1}^{N}Z_{i}$. For hard-sphere packings, however, interparticle overlaps are non-existent, and the above definition of $Z_{i}$ gives $Z=0$. Many authors have instead employed the $C$-dependent contact criterion discussed in Section III.1, i.e. $Z_{i}=\sum_{j\neq i}\Theta[(1+C)\sigma_{ij}-r_{ij}]$, but this approach gives strongly-$C$-dependent $Z$ for small $C$ owing to the divergence of the scaled pair correlation function $g(r_{ij}/\sigma_{ij})$ as $r_{ij}$ approaches $\sigma_{ij}$ from above O’Hern et al. (2003); Donev et al. (2005); Ozawa et al. (2017); Anzivino et al. (2023).
One can mitigate this issue by choosing $C$ values that lie within the well-known plateau in $Z(C)$ Donev et al. (2005), but this complicates rattler identification. For hard-sphere packings, most previous studies have identified rattlers as particles with $Z_{i}<4$ non . Many well-known results for marginally-jammed packings, e.g. the characteristic isostaticity criterion $Z_{\rm J}=6$ and scaling laws such as $Z(\phi)-Z_{\rm J}\sim\sqrt{\phi-\phi_{\rm J}}$, were obtained in systems where rattlers were recursively removed prior to calculating $Z$ O’Hern et al. (2003); Zaccone and Scossa-Romano (2011); Ozawa et al. (2017); Goodrich et al. (2016). Unfortunately, as illustrated in Figure 9, this procedure couples to the $C$-dependence of $Z$.
Figure 9: Protocol dependence of average coordination numbers and rattler fractions for $N=1000$ and various values of $C$ indicated by the color legend. Panel a (b) shows results for systems prepared without (with) SWAP and TDOF moves. Solid (dashed) curves show results for the average coordination number calculated after (before) recursively removing particles with $Z_{i}<4$.
Consistent with the notion that employing SWAP moves allows the $\Gamma\to 0$ limit to be accessed, the differences between the results shown in panels (a-b) are analogous to the differences between results for high and low $\Gamma$ discussed in Ref. Donev et al. (2005). Here we chose $C=10^{-3}$ to perform the analyses described in Section III because it yields $Z_{\rm J}(\Delta)\simeq 6$ for all $\Delta\lesssim 0.25$ in systems prepared without SWAP and TDOF moves, after particles with $Z_{i}<4$ are recursively removed. This allows our results to be straightforwardly compared with the numerous previous studies that have found $Z_{\rm J}=6$.